How do you graph the equation: y= -x + 7? Also how can you tell if it is a function?
Found 2 solutions by richwmiller, rfer:
Answer by richwmiller(17219)

All you need are two points to graph a line if you do them both correctly.
when x=0 y=7
when y=0 x=7
Since it is a line, there is only one y for each x. Therefore it is a function.

Answer by rfer(16322)

This is the form of slope intercept equation
y=mx+b
y=-x+7
m=slope=-1
b=y int=7 (0,7)
(1,6) (2,5)
Does that help?